OpenStudy (anonymous):

Find the general solution of 2cos^2(x) =1

OpenStudy (whpalmer4):

\[2\cos^2 (x) = 1\]\[\cos^2(x) = \frac{1}{2}\]Look at your unit circle — where is that true? Do you see a pattern? Don't forget to figure out where it is true for negative values of \(x\), too.

OpenStudy (anonymous):

But how do you get it in this form? \[\theta=360n \pm \alpha\]

OpenStudy (anonymous):

where n is an integer and \[\cos \theta = \cos\]

OpenStudy (whpalmer4):

first, why don't you tell me the 4 spots on the unit circle that satisfy the equation...

OpenStudy (anonymous):

okay well x= 45, 135, 225, 315 is that correct?

OpenStudy (whpalmer4):



OpenStudy (whpalmer4):

I would normally do this in terms of radians, where the answers would be \[\pi/4, -\pi/4, 3\pi/4, -3\pi/4\] Once we find a formula, we can always convert the amounts to degrees if that's what you really need

OpenStudy (whpalmer4):

looks to me like that is \[\pm(\frac{\pi}{4} + 2\pi n), n\in \text{Integers} \] and \[\pm(\frac{3\pi}{4}+2\pi n), n\in\text{Integers}\]

OpenStudy (anonymous):

I haven't learnt radians though, is there a simpler method?

OpenStudy (whpalmer4):

or just \[\frac{\pi}{4} \pm \frac{\pi}{2}n\] \[\frac{\pi}{4} * \frac{360^\circ}{2\pi} = 45^\circ\]\[\frac{\pi}{2} * \frac{360^\circ}{2\pi} = 90^\circ\] So your angles are just \(45^\circ \pm 90^\circ n\) where \(n\) is an integer and here's a table of the values from n = -10 to 10 \[\begin{array}{cc} x & \cos^2(x)\\\hline\\ -855 {}^{\circ} & \frac{1}{2} \\ -765 {}^{\circ} & \frac{1}{2} \\ -675 {}^{\circ} & \frac{1}{2} \\ -585 {}^{\circ} & \frac{1}{2} \\ -495 {}^{\circ} & \frac{1}{2} \\ -405 {}^{\circ} & \frac{1}{2} \\ -315 {}^{\circ} & \frac{1}{2} \\ -225 {}^{\circ} & \frac{1}{2} \\ -135 {}^{\circ} & \frac{1}{2} \\ -45 {}^{\circ} & \frac{1}{2} \\ 45 {}^{\circ} & \frac{1}{2} \\ 135 {}^{\circ} & \frac{1}{2} \\ 225 {}^{\circ} & \frac{1}{2} \\ 315 {}^{\circ} & \frac{1}{2} \\ 405 {}^{\circ} & \frac{1}{2} \\ 495 {}^{\circ} & \frac{1}{2} \\ 585 {}^{\circ} & \frac{1}{2} \\ 675 {}^{\circ} & \frac{1}{2} \\ 765 {}^{\circ} & \frac{1}{2} \\ 855 {}^{\circ} & \frac{1}{2} \\ 945 {}^{\circ} & \frac{1}{2} \\ \end{array}\]

OpenStudy (whpalmer4):

it's late, I'm not explaining things very well, time to call it a night!

OpenStudy (shiraz14):

Hi, Just to continue where @whpalmer4 has left off (although I think that @whpalmer4 has done a pretty good job here though). From @whpalmer4 's previous post: \[\cos ^{2}x = \frac{ 1 }{ 2 } \] \[\cos x = \pm \sqrt{\frac{ 1 }{ 2 }}\] x = arccos ( \sqrt{\frac{ 1 }{ 2 }}\ ) = 45 deg General solution for x: \[\theta = 360n \pm 45\], where \[n \in \mathbb{Z}^* \].

OpenStudy (anonymous):

Yes, I believe that the general solution for x is \[\theta =360\pm45\] but apparently the textbook answer is \[\theta =180n \pm 45\] So i am quite confused.

OpenStudy (shiraz14):

@fishy13 : Oh yes, I forgot that cosx = +/- sqrt (1/2), which means that x can lie in all 4 quads. So the correct answer should be \[\theta = 180n \pm 45\].

OpenStudy (whpalmer4):

And actually \[\theta = 90n\pm 45^\circ\]is equivalent, although it supplies each point twice...

OpenStudy (whpalmer4):

In any case, @shiraz14, thanks for picking up where I left off! Feel free to do so anytime you see the need...

OpenStudy (shiraz14):

@whpalmer4 : You're welcome. @fishy13 : Yes, this is a special case where you can also have 90n +/- 45 degrees (as @whpalmer4 has stated) simply because 45 degrees is the bisector of 90 degrees. But be careful that this does not always apply, since if theta = 30 degrees, etc, only the conventional argument will work.